# Optimal market-neutral currency trading on the cryptocurrency platform

### v2

@@ 3.2.1 Multivariate Pairs
We want to give a definition and a consistent notation for Pairs. The assets in this research are fiat currencies against ETH such as ETH/USD, ETH/CAD, ETH/GBP. When a pair between ETH/USD and ETH/CAD is formed, the pair will be called as ETH/USD:ETH/CAD, which can be abbreviated to USD:CAD. Theoretically, we can maximum form $k(k-1)/2$ pairs from $k$ assets.
@@ 3.2.2 Buy/Sell Signaling
Arbitrage relies on anomalies. We extract anomalies by comparing normalized spread with standard deviation. We have Open_Threshold and Close_Threshold, the two selected parameters for tuning the trading signals. For a two-asset trading scenario: The trading process between the assets $i$ and $j$ can be generated as follow:
  • Step 1: Logarithmize the Price of $i$ and $j$ to reflect the exponential price growth in the cryptocurrency market.
  • Step 2: Difference logarithmized Price as the Spread array.
  • Step 3: Extract the Standard Deviation from Spread, and normalize the spread array by z-score.
  • Open: If there is not a position opened already, and z-score $>$ Open_Threshold $\times$ Standard Deviation, then open a new position.
  • Close: If there is an open position, and z-score $<$ Close_Threshold $\times$ Standard deviation, then close the position.
  • Step 4: Repeat Open and Close steps until the end of arbitrage.
We give a graphical example to explain the workings of the proposed approach. Table 1 gives the statistics and graphs for a three-asset scenario. Since each asset moves independently, we can see that each pair has a different standard deviation after normalization. The spreads are volatile, and we draw thresholds in auxiliary lines. Every touch of the Open_Threshold represents a position opening signal. On the contrary, if an open position exists between the spread and the spread touches the Close_Threshold, it sends out a position closing signal.
@@
  • In a cut-off period on 17 Jan 2021, a visualized spread trading signalling based on the OPEN_THRESHOLD as 7 and CLOSE_THRESHOLD as 6.
@@ 3.2.3 Position Movements
Suppose we get back from the spreads to the actual asset granularity (Figure 4). The initial principal is distributed amongst the fiat currencies we are trading. Traditionally, we long the loser and short the winner for every positive signal. However, under multiple spread conditions, the trading signals are calculated collectively at every trading interval. The subplot 4(a) displays the ETH price after normalization as price index to express a comparable price growth of ETH in USD, CAD and GBP, and the subplot 4(b) tracks the principal movement of our fiat currencies.
@@
We can see that at time $t_{1}$, GBP starts a rapid decline. Along with that, both relevant spreads GBP:CAD and GBP:USD are shrinking. This is the time ($t_{1}$) for trading signals from spread GBP:CAD and GBP:USD. Our corresponding actions are longing ETH with GBP, and shorting the ETH to USD and CAD. At the latter time $t_{2}$, there is a rebound in ETH price against all currencies. Though all fiat currencies are going up, the USD reversal is weaker than the other two, which makes USD:CAD and USD:GBP spread greater than the open threshold. It triggers the USD to be spent for buying ETH and selling the ETH for CAD and GBP.
By design, it is possible that a fiat currency needs to be longed and shorted at the same time. Besides, the classical scenario of two-asset trading still exists because each spread is monitored individually. An example if the movement that happened at $t_{3}$ does not trigger trading in CAD. The latency of rebound in USD makes only the $z$-score of USD:GBP greater than open thresholds. Therefore the position only opened for the spread USD:GBP. Later after $t_{4}$ when the spread widens, the position opened at $t_{3}$ is closed.
There are more benefits of multivariate asset inclusion. On the one hand, if a fiat currency is receiving a long signal from one pair and the same amount of another short signal, the action can be canceled off on paper to save fees. On the other hand, when one fiat currency reveals a strong anomalous movement, all relevant pairs shall display signals for the principal flowing in/out from the fiat currency. If there is an entry signal in USD:GBP, all the USD can be used to conduct the operation. There will be no USD sitting on the bench and missing out. Though we like to see the perfect mitigation saves on the transaction fee. Signals that can cancel off on each other are rare cases. More often, we have multiple short and long signals that happen simultaneously. And it may receive one short signal from a position closure and another short signal from a new position opening. Therefore, we need to devise a mechanism to determine how exactly the funds should be distributed.
@@ 3.3 An Optimal Approach to Capital Allocation
Resource distribution is a question about how much to invest in each asset every time when there is a trading opportunity. Therefore, at every trading interval, as long as a trading opportunity is detected, the optimization process is executed for a quantitative decision on the amount to invest. We don’t need optimization if we only ever trade a two-asset pair since all the capital (or a pre-determined amount) can be allocated for the long and short positions to be taken. When we have more than two assets, the multilateral relationships require that we ensure the resources distributed on each signal result in the best outcome. The optimization process aims to maximize return with minimum risk under restrictions. In this research, a bi-objective optimization formulation inspired from the Markowitz portfolio allocation (Markowitz 1952) operates to find the optimal investment solution.
@@ 3.3.2 Parameters
    Let $\mathbf{N}$ the a vector of Pairs formed from Currencies, where $|\mathbf{N}|\in\mathbb{Z}^{+}$ represents the number of pairs.
@@
    Let $\mathbf{t}\in[1,2,3,...,T]$ represent the timestamp of execution, and $\mathbf{T}$ denotes the total length of the arbitrage period.
@@
    $\begin{cases}\text{ETH/USD}:[P_{USD,t=1},P_{USD,t=2}...,P_{USD,t=T}]\\ \text{ETH/CAD}:[P_{CAD,t=1},P_{CAD,t=2}...,P_{CAD,t=T}]\\ \text{ETH/GBP}:[P_{USD,t=1},P_{USD,t=2}...,P_{USD,t=T}]\\ \end{cases}$
@@
    Let $\mathbf{r}\in\mathbb{R}^{|C|}$ recording the average returns of ETH price in $|C|$ fiat currencies throughout the arbitrage period and let Expected Profit $\mathbf{EP}\in\mathbb{R}^{|N|\times 2}$ be a matrix for $N$ pairs recording how much profit is expected on average once an arbitrage position is opened. Because each spread consists of two different assets, the speeds of mean-reversion should differ from pair to pair. Hence, we measure a parameter vector $\mathbf{mr}$ of type $\mathbb{R}^{|N|}$ to record the average reversion time to the mean from historical data. Expected Profit is calculated based on the combination of Returns and Mean-reversion speed as follows:
@@
    Let Covariance $\mathbf{cov}\in\mathbb{R}^{N\times 2\times 2}$ be a three dimensional matrix. It contains $2\times 2$ matrices for $N$ pairs. The covariance matrix is used to compute the risk of opening this position. An individual $2\times 2$ matrix from the three-dimensional slice is a variance-covariance matrix for the two fiat currencies in a pair. An example of the $2\times 2$ looks like this below:
@@
    Let Transaction Cost $\mathbf{tc}\in\mathbb{R}^{\geq 0}$ be a constant, which represents the percentage to be deducted from each investment amount as commission.
@@
    Let Trading Weights $\mathbf{TW}\in\mathbb{R}^{|C|}$ be a vector of currency weight occupation for every currency. Since our arbitrage is a continuous process, every trade decision carries historical information. A strong buy signal is guiding the optimization process to invest all the USD into ETH. Because we do not use margin trading, hence the maximum we can invest is 100%. However, if 40% USD has been used to long USD:CAD in an unclosed open position, the maximum we can invest for the next signal falls down to 60%. In this example, the 40% is carried from a weight that is being traded. That is why we give it the name: Trading Weights (TW), and it can be positive or negative.
@@ 3.3.3 Objective
The objective of the optimization is to maximize the return while minimizing the risk. We receive trading signals from spreads, so the decisions for investment weights are in pairs. When we need to long USD and short CAD in the pair USD:CAD, we are effectively applying the weight to action on betting the rise of $r_{USD}$ and the drop of $r_{CAD}$. The expected profit is weights multiply the expected profit presented as:
@@
Where $\odot$ is the Hadarmard product, as the long weights are positive and short weights are negative. We also expect to minimize risk, which means the portfolio has to have as little volatility as possible. According to the modern portfolio theor (Markowitz 1952), the portfolio volatility is $\boldsymbol{w}\times\boldsymbol{cov_{2\times 2}}\times\boldsymbol{w}^{{}^{\prime}}$ for any pair spread in $N$. Once again, as the long weights are positive and short weights are negative, we need to take the Hadamard product with $\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}$ to correctly capture the portfolio covariance in the covariance-variance matrix. Overall, for $N$ pairs we get Equation (2).
@@ 3.3.4 Weight Constraints
$0\leq w_{n,long}\leq 1,\;-1\leq w_{n,short}\leq 0,\;\forall n\in N\$ (3)
Another constraint comes from the weight at the currency level. Every currency should not be spent more than 100% at any time of the trade. Every time the optimization process runs, the process calculates the summation of current trading weight, long weight and short weight to ensure that it does not exceed 100% for each individual currency.
@@ 3.3.5 Market Neutrality
Market Neutrality is the ability to generate stable income regardless of the market condition, and it is one of the major incentives for investors to select pair trading from other quantitative trading methods. To achieve neutrality, we have to ensure that the equivalent amount of ETH in long and short positions are always equal. In that, we need to translate the weights back to the amount of ETH. The price of currency $c$ at time $t$ is attainable as $P_{c,t}$.
In our case, whenever an ETH is bought in one currency, that ETH should be sold instantaneously. Hence, we expect the last constraint is that every pair position reflecting back on assets is equalized in the unit of ETH. Transaction cost tweaks the balance slightly. The amount of ETH received from longing is reduced because of transaction cost. However, the amount of ETH required to get a certain amount of fiat currency increases. The market neutrality constraint is presented in Equation (3.3.5):
@@ 3.3.6 Optimization Equation
Given the above, for the purpose of maximizing the profitability and minimizing the risk, one can optimize the allocation of the principal capital by solving the bi-objective optimization problem in Equation (6). In the optimization process, the ultimate goal is to calculate a set of quantitative weights when opening any position on the pair spreads.
We understand that market participants have different tendencies toward risk aversion. Therefore, we use a coefficient $\lambda$ (default to 1) to measure the degree of risk aversion in the objective. The greater the $\lambda$, the greater the risk aversion. In extreme cases, $\lambda=0$ stands for risk neutrality. The $\lambda$ and portfolio variance form a punishment factor in the objective to cater for investors’ risk preferences.
@@ 4 Simulation
This section gives the process and structure of our simulation (Algorithm 1) on historical data. The process contains multiple stages — screening, hist_analysis, trading and optimisation.
@@
/* Historical Pricing data for ETH against USD, EUR, CAD, GBP */
@@
       /* calculate spread data from assets trading period pricing data */
8       for time in [open_data, close_date] do
@@
                   /* Close position for a pair when there is a position opening and the spread is smaller than the closing threshold */
@@
       /* Calculate standard deviation, expected risk/return among spread */
@@
       /* Set trading threshold between spreads */
@@
The screening (c.f. Section 3.1) determines what assets should be included within our trading bucket. ETH pricing data is extracted from Crypto exchanges against fiat currencies as the input (Algorithm 1 Line 1). We choose which period we trade upon by splitting the pricing data into formation data and trading data. In the formation period, ETH prices against fiat currencies are screened for filling the bucket of fiat currencies with strong correlation and high cointegration test passing rate. Once we dropped the fiat currencies with weak correlation/cointegration relationships, the screened currencies are included in $participating\_assets$. Then we call function hist_analysis.
In the hist_analysis process (Algorithm 1 Line 1), standard deviation, expected return, and expected risk are computed using the historical formation data. Parameters OPEN_THRESHOLD and CLOSE_THRESHOLD are computed using the formation data (c.d. Section 3.2.2). Knowing when to trade is not sufficient enough, we also need to calculate how much to bid/ask on each order. The trading function (Algorithm 1 Line 1) simulates by amount. The first step is to set up $orig\_amount$, which is how much for participating fiat currency we have at the beginning of the trade. Secondly, every two currencies are combined into spreads. Next, we loop through all the trading times and constantly check if any open position meets the closing criteria:
  • if yes, we close the open position to harvest profit, if any. Trading weight vector $TW$ is also updated as described in Section 3.3.
  • if no, we leave it as is and check again in the next time sample
Same as the closing, all the spreads that do not have an ongoing open position need to be checked if an open signal hits at every time sample. Different from closing a position, open-position actions must be considered for how much $orig\_amount$ to spend in each open signal.
To calculate the amount to invest, we call the optimisation function (Algorithm 1 Line 1) by feeding trading data and properties of trading data such as expected return and expected risk to calculate the weights. The optimization is based on Equation 6. Once the optimal weights are calculated, it is returned to the trading function for the continuation of position opening till the next call.
5 Experimental
In this section, we describe the experiments carried out on historical datasets. We first describe the experimental dataset and setup followed by the result of applying the proposed Optimal Trading Technique (OTT)(c.f. Algorithm 1) in different market conditions.

### v3

@@ 3.2.1 Multivariate Pairs
We define and provide a consistent notation for Pairs. ETH/USD or ETH/CAD are the notations of cryptocurrency against fiat currencies. A pair between ETH/USD and ETH/CAD, is referred to as ETH/USD:ETH/CAD, abbreviated as USD:CAD. Theoretically, we can form a maximum of $k(k-1)/2$ pairs from $k$ assets.
@@ 3.2.2 Buy/Sell Signaling
Arbitrage relies on anomalies. We extract anomalies by comparing the normalized spread with the standard deviation. We use two parameters, Open_Threshold and Close_Threshold, to tune the trading signals. For a two-asset trading scenario, the trading process between assets $i$ and $j$ is as follows:
  • Calculate the difference of the logarithmized Price to obtain the Spread array.
  • Normalize the spread array using the Z-Score.
  • Open a new position if there is no existing position and the Z-score $>$ Open_Threshold $\times$ Standard Deviation.
  • Close the position if there is an open position and the Z-score $<$ Close_Threshold $\times$ Standard Deviation.
  • Repeat until the end of the arbitrage period.
@@
  • In a cut-off period on 17-Jan-2021, a visualized spread trading signaling based on the open_threshold as 7 and close_threshold as 6.
We provide a graphical example to explain the proposed approach. Table 1 gives the statistics and graphs for a three-asset scenario. Since each asset moves independently, each pair has a different standard deviation after normalization. The spreads are volatile, and thresholds are represented by auxiliary lines. Every touch of the open_threshold indicates a position opening signal. Conversely, if an open position exists and the spread touches the close_threshold, it signals a position closing.
@@ 3.2.3 Position Movements
Suppose we get back from the spreads to the actual asset granularity (Figure 4). The initial principal is distributed among the fiat currencies we are trading. Traditionally, we long the loser and short the winner for every positive signal. However, multiple spreads might produce conflicting signals. An example is given in Subplot 4(a), which displays the ETH price after normalization as a price index to express a comparable price growth of ETH in USD, CAD, and GBP, while Subplot 4(b) tracks the principal movement of our fiat currencies.
@@
At time $t_{1}$, GBP starts a rapid decline. Both relevant spreads GBP:CAD and GBP:USD shrink, triggering trading signals. Our corresponding actions are longing ETH with GBP and shorting ETH to USD and CAD. At a later time $t_{2}$, there is a rebound in ETH price against all currencies. Although all fiat currencies are rising, the USD reversal is weaker than the others, making the USD:CAD and USD:GBP spreads greater than the open threshold. This triggers the use of USD to buy ETH and sell ETH for CAD and GBP.
By design, it is possible for a fiat currency to be longed and shorted simultaneously. Additionally, the classical scenario of two-asset trading still exists because each spread is monitored individually. For example, the movement at $t_{3}$ does not trigger trading in CAD. The latency of rebound in USD makes only the Z-Score of USD:GBP greater than the open threshold, so a position is opened for the spread USD:GBP. Later, after $t_{4}$, when the spread widens, the position opened at $t_{3}$ is closed.
There are more benefits to including multivariate assets. If a fiat currency receives a long signal from one pair and an equivalent short signal from another, the actions can cancel out on paper, saving fees. Furthermore, when one fiat currency reveals a strong anomalous movement, all relevant pairs will display signals for the principal flowing in or out of that fiat currency. If there is an entry signal in USD:GBP, all the USD can be used to conduct the operation, ensuring no USD sits idle. Though ideally, this mitigation saves on transaction fees, signals that cancel each other are rare. More often, we have multiple short and long signals happening simultaneously. Additionally, one short signal may come from a position closure and another from a new position opening. Therefore, we need to devise a mechanism to determine how funds should be distributed precisely.
@@ 3.3 An Optimal Approach to Capital Allocation
Resource distribution addresses how much to invest in each asset whenever a trading opportunity arises. At every trading interval, as long as a trading opportunity is detected, the optimization process is executed to make a quantitative decision on the amount to invest. Optimization is unnecessary if we only trade a two-asset pair since all the capital (or a pre-determined amount) can be allocated for the long and short positions. However, with more than two assets, the multilateral relationships necessitate ensuring that the resources distributed on each signal result in the best outcome. The optimization process aims to maximize return with minimum risk under given constraints. In this research, a bi-objective optimization formulation inspired by the Markowitz portfolio allocation (Markowitz 1952) is used to find the optimal investment solution.
@@ 3.3.2 Parameters
    Let $\mathbf{N}$ be a vector of Pairs formed from Currencies, where $|\mathbf{N}|\in\mathbb{Z}^{+}$ represents the number of pairs.
@@
    Let $\mathbf{t}\in[1,2,3,...,T]$ represent the timestamp of execution, and $T$ denotes the total length of the arbitrage period.
@@
    $\begin{cases}\text{ETH/USD}:[P_{USD,t_{1}},P_{USD,t_{2}}...,P_{USD,t_{T}}]\\ \text{ETH/CAD}:[P_{CAD,t_{1}},P_{CAD,t_{2}}...,P_{CAD,t_{T}}]\\ \text{ETH/GBP}:[P_{GBP,t_{1}},P_{GBP,t_{2}}...,P_{GBP,t_{T}}]\\ \end{cases}$
@@
    Let $\mathbf{r}\in\mathbb{R}^{|C|}$ record the average returns of ETH price in $|C|$ fiat currencies throughout the arbitrage period, and let Expected Profit $\mathbf{EP}\in\mathbb{R}^{|N|\times 2}$ be a matrix for $N$ pairs recording the average expected profit once an arbitrage position is opened. Since each spread consists of two different assets, the speeds of mean-reversion should differ from pair to pair. Therefore, we measure a parameter vector $\mathbf{mr}$ of type $\mathbb{R}^{|N|}$ to record the average reversion time to the mean from historical data. Expected Profit is calculated based on the combination of Returns and Mean-reversion speed as follows:
@@
    Let Covariance $\mathbf{cov}\in\mathbb{R}^{N\times 2\times 2}$ be a three-dimensional matrix. It contains $2\times 2$ matrices for $N$ pairs. The covariance matrix is used to compute the risk of opening a position. An individual $2\times 2$ matrix from the three-dimensional slice is a variance-covariance matrix for the two fiat currencies in a pair. An example of the $2\times 2$ matrix is shown below:
@@
    Let Transaction Cost $\mathbf{tc}\in\mathbb{R}^{\geq 0}$ be a constant, representing the percentage deducted from each investment amount as commission.
@@
    Let Trading Weights $\mathbf{TW}\in\mathbb{R}^{|C|}$ be a vector of currency weight occupation for every currency. Since our arbitrage is a continuous process, every trade decision carries historical information. A strong buy signal guides the optimization process to invest all the USD into ETH. Because we do not use margin trading, the maximum we can invest is 100%. However, if 40% USD has been used to long USD:CAD in an unclosed open position, the maximum we can invest for the next signal falls to 60%. In this example, the 40% is carried from a weight that is being traded. That is why we give it the name: Trading Weights (TW), and it can be positive or negative.
@@ 3.3.3 Objective
The objective of the optimization is to maximize the return while minimizing the risk. We receive trading signals from spreads, so the decisions for investment weights are in pairs. When we need to long USD and short CAD in the pair USD:CAD, we are effectively applying the weight to bet on the rise of $r_{USD}$ and the drop of $r_{CAD}$. The expected profit is calculated as the weights multiplied by the expected profit, presented as:
@@
where $\odot$ is the Hadamard product, with long weights positive and short weights negative. We also aim to minimize risk, meaning the portfolio should have as little volatility as possible. According to modern portfolio theory (Markowitz 1952), the portfolio volatility is $\boldsymbol{w}\times\boldsymbol{cov_{2\times 2}}\times\boldsymbol{w}^{{}^{\prime}}$ for any pair spread in $N$. Since long weights are positive and short weights are negative, we need to take the Hadamard product with $\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}$ to correctly capture the portfolio covariance in the covariance-variance matrix. Overall, for $N$ pairs, we get Equation (2).
@@ 3.3.4 Weight Constraints
$0\leq w_{n,long}\leq 1,\;-1\leq w_{n,short}\leq 0,\;\forall n\in N$ (3)
Another constraint comes from the weight at the currency level. Every currency should not be spent more than 100% at any time of the trade. Every time the optimization process runs, it calculates the summation of the current trading weight, long weight, and short weight to ensure that it does not exceed 100% for each individual currency.
@@ 3.3.5 Market Neutrality
Market Neutrality is the ability to generate stable income regardless of the market condition, and it is one of the major incentives for investors to select pair trading from other quantitative trading methods. Market neutrality requires an equivalent distribution of investment in the long and short positions. In that, we need to translate the weights back to the amount of ETH. The price of currency $c$ at time $t$ is attainable as $P_{c,t}$.
In our case, whenever an ETH is bought in one currency, that ETH should be sold instantaneously. Hence, we expect the last constraint is that every pair position reflecting back on assets is equalized in the unit of ETH. Transaction cost tweaks the balance slightly. The amount of ETH received from longing is reduced because of transaction costs. However, the amount of ETH required to get a certain amount of fiat currency increases. The market neutrality constraint is presented in Equation (3.3.5):
@@ 3.3.6 Optimization Equation
Given the above, for the purpose of maximizing profitability and minimizing risk, one can optimize the allocation of the principal capital by solving the bi-objective optimization problem in Equation (6). The ultimate goal of the optimization process is to calculate a set of quantitative weights when opening any position on the pair spreads.
We understand that market participants have different tendencies toward risk aversion. Therefore, we use a coefficient $\lambda$ (default to 1) to measure the degree of risk aversion in the objective. The greater the $\lambda$, the greater the risk aversion. In extreme cases, $\lambda=0$ stands for risk neutrality. The $\lambda$ and portfolio variance form a penalty factor in the objective to cater to investors’ risk preferences.
@@ 4 Simulation
This section describes the process and structure of our simulation (Algorithm 1) on historical data. The process consists of multiple stages: screening, hist_analysis, trading, and optimisation.
@@
/* Historical pricing data for cryptocurrency against fiat currencies */
@@
       /* Calculate spread data from asset trading period pricing data */
8       for time in [open_date, close_date] do
@@
                   /* Close position for a pair when there is an open position and the spread is smaller than the closing threshold */
@@
       /* Calculate standard deviation, expected risk/return among spreads */
@@
       /* Set trading thresholds between spreads */
@@
The screening (c.f. Section 3.1) determines which assets should be included within our trading bucket. ETH pricing data is extracted from crypto exchanges against fiat currencies as the input (Algorithm 1 Line 1). We choose the trading period by splitting the pricing data into formation data and trading data. In the formation period, ETH prices against fiat currencies are screened to fill the bucket with fiat currencies that have strong correlation and high cointegration test passing rates. Once we drop the fiat currencies with weak correlation/cointegration relationships, the screened currencies are included in participating_assets. Then we call the function hist_analysis.
In the hist_analysis process (Algorithm 1 Line 1), standard deviation, expected return, and expected risk are computed using the historical formation data. Parameters open_threshold and close_threshold are computed using the formation data (c.f. Section 3.2.2). Knowing when to trade is not sufficient; we also need to calculate how much to bid/ask on each order. The trading function (Algorithm 1 Line 1) simulates by amount. The first step is to set up $orig\_amount$, which is the initial amount of each participating fiat currency at the beginning of the trade. Secondly, every two currencies are combined into spreads. Next, we loop through all the trading times and constantly check if any open position meets the closing criteria:
  • If yes, we close the open position to harvest profit. The trading weight vector $TW$ is also updated as described in Section 3.3.
  • If no, we leave it as is and check again in the next time sample.
Similarly, all the spreads that do not have an ongoing open position need to be checked at every time sample to see if an open signal is triggered. Different from closing a position, open-position actions must consider how much $orig\_amount$ to spend on each open signal.
To calculate the amount to invest, we call the optimisation function (Algorithm 1 Line 1) by feeding trading data and properties of trading data such as expected return and expected risk to calculate the weights. The optimization is based on Equation 6. Once the optimal weights are calculated, they are returned to the trading function for the continuation of position opening until the next call. The code implementing our methodology is available in a public repository [^1] The source code can be accessed at https://github.com/Hongshen-Yang/OTT.
5 Experiment
In this section, we apply our technique on historical ETH datasets against USD, EUR, CAD, and GBP. We select three market conditions to test the applicability of Optimal Trading Technique (c.f. Algorithm 1), including a bull-run market (01-Jan-2021 to 01-Jan-2022, 365 days), a bear-run market (01-Jan-2022 to 01-Jan-2023, 365 days), and a full-cycle market (15-Jan-2021 to 01-Oct-2022, 625 days) for 1min, 5min and 60min.